# VBench-2.0: Advancing Video Generation Benchmark Suite for Intrinsic Faithfulness

### v1

@@ H More Details on Prompt Suite
HunyuanVideo [72] 50.64% / 51.37% 52.75% / 55.07% 51.46% / 56.35% 49.72% / 52.45% 52.78% / 60.04% 45.74% / 30.46% 46.60% / 49.28% 49.43% / 52.68% 49.61% / 52.36%
CogVideoX-1.5 [86] 49.36% / 48.07% 53.72% / 54.70% 51.68% / 52.13% 53.28% / 51.34% 55.74% / 54.46% 52.50% / 65.37% 46.19% / 44.08% 49.04% / 42.24% 41.91% / 41.32%
Sora [53] 49.68% / 49.60% 42.98% / 43.16% 43.60% / 30.81% 43.94% / 43.49% 40.56% / 35.87% 50.74% / 49.35% 42.08% / 42.64% 49.43% / 44.83% 55.17% / 52.94%
Kling 1.6 [69] 50.32% / 50.97% 50.55% / 47.07% 53.25% / 60.73% 53.06% / 52.73% 50.93% / 49.63% 51.02% / 54.81% 65.12% / 63.99% 52.11% / 60.25% 53.31% / 53.37%
Correlation 97.36% 90.95% 98.73% 89.93% 91.17% 97.65% 97.95% 87.97% 93.70%
I Human Preference Annotation Details
I.1 Data Preparation
There are two types of annotation formats in VBench-2.0.
I.2 Labeling Instructions.
The annotation process follows VBench but incorporates refinements in interface design and evaluation methodology. Since VBench-2.0 introduces multiple dimensions with detailed multi-question evaluations and long text prompts (some exceeding 150 words), we enhance the interface for improved readability and efficiency. Instead of displaying extensive text in video titles, we sequentially list all key annotation instructions directly within the interface. This structured layout allows annotators to efficiently reference the necessary details while conducting evaluations.
I.3 Win Ratio.
Given human annotations, we calculate the win ratio for each model. During pairwise comparisons, if a model’s video is preferred, it scores 1, while the other model scores 0. In case of a tie, both models score 0.5. The win ratio for each model is computed as the total score divided by the number of pairwise comparisons. We show the result in Table S4
I.4 Quality Assurance
The annotation process maintains the rigorous quality control measures established in VBench while further refining the review criteria. We randomly sample 20% of the annotated pairs for verification, with a required success rate of 95%
To quantify the annotation effort across all 18 evaluation dimensions, we account for the cumulative time spent on documentation drafting, initial trials, formal annotation, and re-annotation. In total, the process required 15.75 hours of individual effort and 284 hours across 18 annotators, reflecting the scale of the task and our commitment to ensuring high-quality human preference annotations. Most evaluation dimensions go through 2 rounds of trial labeling, official labeling, and post-labeling verification
J Video Generation Models in Evaluation
To assess our benchmark against recent advancements, we utilize four models for comparison, with more to be included as they become open-sourced. We introduce the four models in detail and a unified prompt refiner to pre-process the text prompt as follows.
Kling 1.6 [69]. We adopt the standard API version of Kling 1.6. For each prompt, we sample 241 continuous frames of size 720$\times$1280 at 24 frames per second (FPS). For the classifier-free guidance (cfg) scale, we used the official default value of 0.5.
Sora-480p [53]. Sora represents one of the earliest large-scale video generation models to emerge. For each prompt, we directly perform sampling on the official website, containing 150 continuous frames of size 480$\times$854 at 30 FPS (i.e., the low-resolution version of Sora).
HunyuanVideo [72]. HunyuanVideo is a powerful open-sourced video generation model. It is equipped with a pre-trained Multimodal Large Language Model (MLLM) [67] to better understand text prompts, supporting a large maximum token length and incorporating a dual-branch diffusion transformer generative architecture, which produces visually convincing results. We use the best version of HunyuanVideo with no classifier-free guidance. For each prompt, we sample 129 continuous frames of size 720$\times$1280 at 24 FPS. All of the hyper-parameters are followed the default value in the official inference code. The initial random seed is set to 42 during sampling.
CogVideoX-1.5 [86]. CogVideoX-1.5 is a transformer-based video generation model, which is equipped with a 3D causal VAE to compress the spatial and temporal dimensions and an expert transformer generative architecture to achieve better text-video alignment. We use the best version of it (5B) to evaluate. For each prompt, we sample 161 continuous frames of size 768$\times$1360 at 16 FPS. The initial random seed is also set to 42 for a fair comparison.

### v2

@@ IV-A Video Generation Models in Evaluation
HunyuanVideo [31] 50.64% / 51.37% 52.75% / 55.07% 51.63% / 56.35% 49.61% / 52.45% 52.41% / 60.04% 46.85% / 30.46% 46.60% / 49.28% 49.43% / 52.68% 51.36% / 50.88%
CogVideoX-1.5 [59] 49.36% / 48.07% 53.72% / 54.70% 51.40% / 52.13% 53.39% / 51.34% 53.89% / 54.46% 52.31% / 65.37% 46.19% / 44.08% 49.04% / 42.24% 44.74% / 43.27%
Sora [28] 49.68% / 49.60% 42.98% / 43.16% 43.77% / 30.81% 43.83% / 43.49% 42.04% / 35.87% 51.30% / 49.35% 42.08% / 42.64% 49.43% / 44.83% 52.53% / 53.80%
Kling 1.6 [29] 50.32% / 50.97% 50.55% / 47.07% 53.20% / 60.73% 53.17% / 52.73% 51.67% / 49.63% 49.54% / 54.81% 65.12% / 63.99% 52.11% / 60.25% 51.36% / 52.05%
Correlation 97.36% 90.95% 99.31% 89.28% 91.82% 89.59% 97.95% 87.97% 99.11%
IV-B VBench-2.0 Evaluation Results
For each sub-ability dimension, videos were generated using the models based on the corresponding prompt suite described in Section III-B. The evaluation method introduced in Section III-A is then applied to obtain numerical scores between 0 and 1, where a higher value indicate relatively stronger performance in that dimension. The evaluation results of the four video generative models are summarized in Table II and visualized in Figure 2. Additionally, for each of the five major dimensions, we visualize the evaluation results of one model pair to demonstrate the accuracy of our method in Figure 6 (Complex Landscape in Controllability), Figure 5 (Diversity in Creativity), Figure 4 (Human Anatomy in Human Fidelity), Figure 7 (Mechanics in Physics) and Figure 8 (Instance Preservation in Commonsense).
IV-C Human Alignment of VBench-2.0
To ensure that VBench-2.0’s evaluation aligns closely with human judgment across all evaluation dimensions, we conducted human preference labeling on a large set of generated videos, following the approach of VBench [25, 26]. Specifically, we computed the correlation between our evaluation results and human annotations. Figure 11 and Table IV presents the correlation plot and numerical win ratios respectively, illustrating the alignment between human judgment and VBench-2.0 evaluations in terms of model-level win ratios across video generation models in each dimension.
V Insights and Discussions
V-A Characteristics of Recent SOTA Models
V-B Key Limitations of Recent SOTA Models
V-C The Role of Prompt Engineering
In modern video generation, the Prompt Refiner rewrites or augments input text prompts to enhance video generation quality. For implementation details, please refer to the Sec IV-A. Below, we highlight several notable and potentially surprising observations.
Prompting Partially Compensates for Physical Reasoning Gaps. Physics is often challenging in video generation. However, with the exception of Sora-480p, models guided by the external Prompt Refiner demonstrated reasonably strong performance in the Physics dimension. This suggests that even without an inherent understanding of physical laws, models can be steered toward physically plausible outcomes through carefully designed prompts. Therefore, the core difficulty may lie less in physical reasoning itself, and more in achieving precise video-text alignment.
Limited Impact on Knowledge-Driven Dimensions. For dimensions that rely on model’s intrinsic visual understanding and prior knowledge, like Human Fidelity, Camera Motion, Geometry, and Commonsense, we observe no consistent performance trend from models using different Prompt Refiners. These aspects likely extend beyond the scope of logical inference or direct text-to-video mapping, limiting the Prompt Refiner’s influence. This suggests that success in these areas may depend less on prompt engineering and more on underlying data quality and model architecture.
V-D Superficial Faithfulness vs. Intrinsic Faithfulness: Do Not Miss Out on Any Pillar
Superficial Faithfulness (e.g., cinematographic quality) often shapes the first impression viewers get from a video. As a result, models that produce aesthetically pleasing and smooth outputs are frequently perceived as “better”. However, this perception could be misleading. In practice, Intrinsic Faithfulness, which includes elements like storytelling, logical progression, and world simulation, is equally important for determining a model’s potential for real-world applications. For example, as shown in Figure 2 and Table II, CogVideoX performs relatively well across many VBench-2.0 dimensions, though its visual Quality Score in VBench suggests room for improvement compared to models like Sora and HunyuanVideo. Conversely, HunyuanVideo produces visually impressive results, though its performance in many structure-driven dimensions in VBench-2.0 suggests opportunities for further growth in those areas. These observations highlight a common bias toward relying primarily on visual quality when judging video generation models. To address this, we encourage the community to use both VBench and VBench-2.0 together, enabling a more comprehensive and in-depth evaluation across both Superficial Faithfulness and Intrinsic Faithfulness.
V-E Challenge, Current Solution and Future Work of Video Generation Evaluation
Challenge. Since Intrinsic Faithfulness is more difficult to accurately evaluate compared to Superficial Faithfulness ( i.e., evaluation models need to have capabilities that include complex scene understanding, commonsense reasoning, and physical world perception), the demand for using large models such as VLMs and LLMs has gradually increased. However, in practical applications, dimensions such as Human Anatomy and Motion Rationality are areas where current large models cannot serve as reliable evaluators. The main reasons can be summarized into two points: 1) Training data gap: Real-world video data used for training VLMs does not contain anomaly or unnatural entity states that are commonly present in generated videos; 2) Capability limitations: VLMs perform poorly in aspects such as object counting and 3D understanding.
Current Solution. In this paper, to address the above two challenges, we construct the corresponding solutions: 1) For different tasks and scenarios, we build anomaly specialists to collect abnormal data regarding human body, motion, entity splitting and merging. This data is then mixed with real datasets for training, achieving high accuracy while maintaining generalizability. 2) Currently, there are two mainstream approaches: fine-tuning-based and non-fine-tuning methods. Fine-tuning-based methods can easily overfit to the generated training data (due to limited data size), achieving high test accuracy. However, since most generated videos are less than 10 seconds in length, models tuned in this way struggle to generalize to longer videos in the future, and the overall generalization ability of the model is also a major concern. Due to the above drawbacks, VBench-2.0 adopts a non-fine-tuning approach. We break down a large problem into multiple visual phenomena that can be understood by VLMs, and design redundant questions to improve the accuracy of VLMs. For dimensions that require high-level understanding ability, we further integrate LLM into the evaluation.
Future Work. We believe that some dimensions will inevitably require the understanding capabilities of large models to be resolved in the future. However, other dimensions (particularly those involving content that should not exist in real data) indeed need to be handled by specialized expert models, as outlined in the two points mentioned above. Here, following the above classification, we outline the key directions that future next-generation video generation evaluation methods should explore: 1) Although the types of anomalies judged by different dimensions vary (Human Anatomy and Instance Preservation evaluate anomalies in human motion, appearance anomalies and entity state anomalies, respectively), there is an overlap between these anomaly types. For example, during the entity splitting process, limb anomalies will also occur. Therefore, integrating abnormal data to train a unified clip-level anomaly detection model can help the model capture both the differences and commonalities across various anomalies, thereby enabling the development of a more generalized and unified anomaly detection large model. 2) Accuracy and sustainability are two important pillar off a good benchmark. Fine-tuning-based methods cannot meet the sustainability requirement under current circumstances (unless transform into a clip-level formulation that is independent of video length and construct a virtual database with the same scale as the VLM training data). Therefore, what needs to be done in the future is to more thoroughly evaluate existing models’ understanding of real-world physical laws and 3D scenes (e.g., VLM benchmarks for 3D space reasoning [92], cinematography understanding [93] and physics understanding [94, 95]). When a truly capable and qualified intelligent large model emerges in the future, it will mark the beginning of the next generation of video generation benchmarks.
VI Conclusion
While recent video generative models have achieved superficial faithfulness, true progress requires advancing towards intrinsic faithfulness, ensuring adherence to physical laws, commonsense reasoning, and structured coherence. To address this, we introduced VBench-2.0, a benchmark assessing models on five key dimensions beyond superficial faithfulness. To accurately evaluate these more challenging dimensions, we fully explore the capabilities of state-of-the-art vision-language models (VLMs) and large language models (LLMs). We build generalists in dimensions where they excel, and employ various specialists for assessment in areas where they are less proficient. VBench-2.0 complements VBench by expanding evaluation to deeper aspects of video generation, providing a multi-dimensional and human-aligned evaluation framework. We believe that VBench-2.0 is an important contribution to the video generation community, shaping the field into its next era.
Future Work. We will continually add more video generative models to VBench-2.0 when they become available.
Acknowledgments
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).